# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ Appendix B An amalgamation lemma
In this section, we state and prove an amalgamation lemma, which is a special case of a result from an unpublished work by Solecki and Tsankov. We will need this lemma for the construction in Section 5. For the sake of completeness, we provide a direct proof of this particular case here.
@@
  1. $(1)$
    $L$ is a finite linear graph, $M$, $M^{\prime}$ denote finite linear graphs,
  2. $(2)$
    $f:M\to L$, $f^{\prime}:M^{\prime}\to L$ are epimorphisms,
  3. $(3)$
    $_{-}, m+m_{+}$ ($^{\prime}_{-}, m+m^{\prime}_{+}$, resp.) are endpoints of $M$ ($M^{\prime}$, resp.) which satisfy
    ${\rm tp}^{m_{-},M}(f)={\rm tp}^{m^{\prime}_{-},M^{\prime}}(f^{\prime}).$
    ${\rm tp}^{m_{+},M}(f)={\rm tp}^{m^{\prime}_{+},M^{\prime}}(f^{\prime}).$
Then there exist
  • a finite linear graph $O$ with endpoints $_{-}, o+o_{+}$,
  • and epimorphisms $g:O\to M$, $g^{\prime}:O\to M^{\prime}$ with
@@ 3. $(3)$
  • $g(o_{-})=m_{-}$,
  • $g^{\prime}(o_{-})=m^{\prime}_{-}$,
  • $g(o_{+})=m_{+}$,
  • $g^{\prime}(o_{-})=m^{\prime}_{+}$.
Moreover, if we only assume ${\rm tp}^{m_{-},M}(f)={\rm tp}^{m^{\prime}_{-},M^{\prime}}(f^{\prime}),$ then there exists $g:O\to M$, $g^{\prime}:O\to M^{\prime}$ with
$f\circ g=f^{\prime}\circ g^{\prime},$
and an endpoint $o_{-}$ such that $g(o_{-})=m_{-}$, $g^{\prime}(o_{-})=m^{\prime}_{-}$.
@@
Suppose that $L$, $M$, $M^{\prime}$ are finite linear graphs, $f\in{\rm Epi}(M,L)$, $f^{\prime}\in{\rm Epi}(M^{\prime},L)$, and the endpoints $l_{+},l_{-}$ of $L$, $m_{+},m_{-}$ of $M$, $m^{\prime}_{+},m^{\prime}_{-}$ of $M^{\prime}$ satisfy $f(m_{-})=l_{-}$, $f(m_{+})=l_{+}$, $f^{\prime}(m^{\prime}_{-})=l_{-}$, $f^{\prime}(m^{\prime}_{+})=l_{+}$.
Then, for some finite linear graph $O$ and epimorphisms $g:O\to M$, $g^{\prime}:O\to M^{\prime}$ we have $f\circ g=f^{\prime}\circ g^{\prime}$, and one endpoint of $O$ is mapped to $m_{-}$ by $g$ ($m^{\prime}_{-}$ by $g^{\prime}$, resp.), while the other endpoint is mapped to $m_{+}$ and $m^{\prime}_{+}$.
@@
Extend the linear graphs $M$, $M^{\prime}$, $L$ by two points as follows. We let
$M_{ext}=M\cup\{d_{-},d_{+}\},$
such that $d_{-}$ is connected with $m_{-}$, and $d_{+}$ is connected with cm+cm{+}. Similarly,
$M^{\prime}_{\rm ext}=M^{\prime}\cup\{d^{\prime}{-},d^{\prime}_{+}\},$
@@
$L_{\rm ext}=L\cup\{l^{*}_{-},l^{*}_{+}\}.$
It is easy to see that $\tilde{f}:M_{\rm ext}\to L_{\rm ext}$, defined by the identities $\tilde{f}^{\prime}{\restriction}M=f$,
(B.1) $\begin{array}[]{l}\tilde{f}(d_{-})=l^{*}_{-}\\ \tilde{f}(d_{+})=l^{*}_{+}\end{array}$
is an epimorphism. Similarly, letting $\tilde{f}:M\to L_{\rm ext}$ denote the mapping that satisfies $\tilde{f}^{\prime}{\restriction}M^{\prime}=f^{\prime}$,
(B.2) $\begin{array}[]{l}\tilde{f}^{\prime}(d_{-}^{\prime})=l^{*}_{-},\\ \tilde{f}^{\prime}(d^{\prime}_{+})=l^{*}_{+}\end{array}$
@@
Now it is easily checked that recalling (B.1), (B.2) and $l^{*}_{-},l^{*}_{+}\in L_{\rm ext}\setminus L$, for every $a\in O$
$\tilde{f}\circ\tilde{g}(a)=l^{*}_{-}\ \iff\ \tilde{g}(a)=d_{-},$
@@
$\tilde{f}^{\prime}\circ\tilde{g}^{\prime}(a)=l^{*}_{-}\ \iff\ \tilde{g}^{\prime}(a)=d_{-}^{\prime}.$
@@
Again by minimality one easily checks that if $a\in O$ is adjacent to an endpoint of $O$, then
  • either $\tilde{f}\circ\tilde{g}(a)=\tilde{f}^{\prime}\circ\tilde{g}^{\prime}(a)=l_{-}$, and $\tilde{f}(a)=m_{-}$, $\tilde{f}^{\prime}(a)=m^{\prime}_{-}$,
  • or $\tilde{f}\circ\tilde{g}(a)=\tilde{f}^{\prime}\circ\tilde{g}^{\prime}(a)=l_{+}$, and $\tilde{f}(a)=m_{+}$, $\tilde{f}^{\prime}(a)=m^{\prime}_{+}$

### v2

@@ Appendix B An amalgamation lemma
In this section, we state and prove two amalgamation lemmas. The first one of which was known to the second author for a long time and both are special cases of a result from an unpublished work by Solecki and Tsankov. For the sake of completeness, we provide proofs of these particular cases here.
The first lemma will be needed in Section 4(B).
Lemma B.1 (Solecki).
Supposed that $I$, $J$, $J^{\prime}$ be finite linear graphs (objects in $\mathcal{P}$}), $f:J\to I$, $f^{\prime}:J^{\prime}\to I$ are epimorphisms (in $\mathcal{P}$}), and $a_{J}$ and $a_{J^{\prime}}$ are endpoints of $J$ and $J^{\prime}$, respectively. If
${\rm tp}^{a_{J},J}(f)={\rm tp}^{a_{J^{\prime}},J^{\prime}}(f^{\prime}),$
then there exist epimorphisms $g:L\to J$, $g^{\prime}:L\to J^{\prime}$ and an endpoint $a_{L}$ of $L$ with
$f\circ g=f^{\prime}\circ g^{\prime},$
and
$g(a_{L})=a_{J},\;g^{\prime}(a_{L})=a_{J^{\prime}}.$
Lemma B.1 will follow from Lemma B.2 below. We will need Lemma B.2 also for the construction in Section 5.
@@
  1. II, $J$, $J^{\prime}$ are finite linear graphs (objects in $\mathcal{P}$}),
  2. $f:J\to I$, $f^{\prime}:J^{\prime}\to I$ are epimorphisms (in $\mathcal{P}$}),
  3. $a_{J}$, $b_{J}$ ($a_{J^{\prime}}$, $b_{J^{\prime}}$, resp.) are endpoints of $J$ ($J^{\prime}$, respectively) which satisfy
    ${\rm tp}^{a_{J},J}(f)={\rm tp}^{a_{J^{\prime}},J^{\prime}}(f^{\prime})\;\hbox{ and }\;{\rm tp}^{b_{J},J}(f)={\rm tp}^{b_{J^{\prime}},J^{\prime}}(f^{\prime}).$
Then there exist a finite linear graph $L$ with endpoints $a_{L}$, $b_{L}$ and epimorphisms $g:L\to J$, $g^{\prime}:L\to J^{\prime}$ with
@@ 2. $f:J\to I$, $f^{\prime}:J^{\prime}\to I$ are epimorphisms (in $\mathcal{P}$}),
$g(a_{L})=a_{J},\;g^{\prime}(a_{L})=a_{J^{\prime}},\;g(b_{L})=b_{J},\;g^{\prime}(b_{L})=b_{J^{\prime}}.$
@@
Suppose that $I$, $J$, $J^{\prime}$ are finite linear graphs, $f\in{\rm Epi}(J,I)$, $f^{\prime}\in{\rm Epi}(J^{\prime},I)$, and the endpoints $a_{J},b_{J}$ of $J$, $a_{J^{\prime}},b_{J^{\prime}}$ of $J^{\prime}$ and $a_{I},b_{I}$ of $I$ satisfy $f(a_{J})=a_{I}$, $f(b_{J})=b_{I}$, $f^{\prime}(a_{J^{\prime}})=a_{I}$, $f^{\prime}(b_{J^{\prime}})=b_{I}$.
Then, for some finite linear graph $L$ and epimorphisms $g:L\to J$, $g^{\prime}:L\to J^{\prime}$ we have $f\circ g=f^{\prime}\circ g^{\prime}$, and one endpoint of $L$ is mapped to $a_{J}$ by $g$ ($a_{J^{\prime}}$ by $g^{\prime}$, resp.), while the other endpoint is mapped to $b_{J}$ and $b_{J^{\prime}}$.
@@
Extend the linear graphs $J$, $J^{\prime}$, $I$ by two points as follows. We let
$J_{ext}=J\cup\{c_{J},d_{J}\},$
such that $c_{J}$ is connected with $a_{J}$, and $d_{J}$ is connected with $b_{J}$. Similarly,
$J^{\prime}_{\rm ext}=J^{\prime}\cup\{c_{J^{\prime}},d_{J^{\prime}}\},$
@@
$I_{\rm ext}=I\cup\{c_{I},d_{I}\}.$
It is easy to see that $\tilde{f}:J_{\rm ext}\to I_{\rm ext}$, defined by the identities $\tilde{f}{\restriction}J=f$,
(B.1) $\begin{array}[]{l}\tilde{f}(c_{J})=c_{I},\\ \tilde{f}(d_{J})=d_{I}\end{array}$
is an epimorphism. Similarly, letting $\tilde{f}^{\prime}:J^{\prime}_{\rm ext}\to I_{\rm ext}$ denote the mapping that satisfies $\tilde{f}^{\prime}{\restriction}J^{\prime}=f^{\prime}$,
(B.2) $\begin{array}[]{l}\tilde{f}^{\prime}(c_{J^{\prime}})=c_{I},\\ \tilde{f}^{\prime}(d_{J^{\prime}})=d_{I}\end{array}$
@@
Now recalling (B.1), (B.2) and $c_{I},d_{I}\in I_{\rm ext}\setminus I$, it is easily checked that for every $l\in L$
$\tilde{f}\circ\tilde{g}(l)=c_{I}\ \iff\ \tilde{g}(l)=c_{J},$
@@
$\tilde{f}^{\prime}\circ\tilde{g}^{\prime}(l)=c_{I}\ \iff\ \tilde{g}^{\prime}(l)=c_{J^{\prime}}$
@@
(and similarly with $d_{I}$, $d_{J}$, $d_{J^{\prime}}$). Again by minimality one easily checks that if $l\in L$ is adjacent to an endpoint of $L$, then
  • either $\tilde{f}\circ\tilde{g}(l)=\tilde{f}^{\prime}\circ\tilde{g}^{\prime}(l)=a_{I}$, and $\tilde{g}(l)=a_{J}$, $\tilde{f}^{\prime}(l)=a_{J^{\prime}}$,
  • or $\tilde{f}\circ\tilde{g}(l)=\tilde{f}^{\prime}\circ\tilde{g}^{\prime}(l)=b_{I}$, and $\tilde{f}(l)=b_{J}$, $\tilde{f}^{\prime}(l)=b_{J^{\prime}}$.